# Spectral Differential Network Analysis for High-Dimensional Time Series

### v1

@@ A.2 Functional Dependence Framework
A.3 Proof of Theorem 1
Concentration Bound on $\hat{f}_{l}$ Around $f_{l}$.
We begin our proof by establishing a concentration inequality on
$\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}$
for each condition $i$ using Theorem 3.1 from (Fiecas et al., 2019). Under Assumption 2 we have that for any $\delta>0$ and $H>0$ and for $C_{1}>0$ depending only on $\kappa$ and $C_{0}$
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,$
where $C_{2}$ is constant in $n_{l}$ but depends on $C_{1}$ and $H$. Note that this same inequality holds for both conditions $l=1,2$ because Assumption 2 holds for both conditions. We allow $n_{l}$ to differ between conditions for full generality. We can then select $M_{n_{l}}$ to minimize the deviations. That is we select $M_{n_{l}}$ to minimize $\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}$. The algebra is omitted but setting $\frac{\partial{}}{\partial{M_{n_{l}}}}\left[\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right]=0$ and solving we get $M^{*}_{n_{l}}=\left(\frac{8}{C_{1}}\left(\frac{1}{2}+\delta\right)\right)^{2/(3+2\delta)}n_{l}^{\frac{2+2\delta}{3+2\delta}}$. It can be shown that the second derivative is positive on $\delta>0$ so this is indeed a global minimum (on $\delta>0$). Plugging this back into the equation yields
@@ A.3 Proof of Theorem 1
Letting $\delta\rightarrow 0$ the RHS of our concentration bound becomes $C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}$ and we get for any $H>0$ and for an optimal bandwidth choice of $M_{n_{l}}^{*}=O(n_{l}^{2/3})$ we get
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,.$ (5)
Controlling Deviations of Sample Loss Around True Loss.
Throughout our theoretical analysis we will need to control the deviations of the second derivative of our sample loss around the true value. That is we will need to control
$\left\|0.5\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}\right)-0.5\left(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1}\right)\right\|_{\infty}\,.$
Note that both
@@
First, note that $\|\Sigma_{l}\|_{\infty}\leq\|f_{l}\|_{\infty}$ and denote $C_{f}=\max(\|f_{1}\|_{\infty},\|f_{2}\|_{\infty})$. Using the optimal error bound on the smoothed periodograms from (5) we get that with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
@@
This also holds for $\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty}$. If we define
@@
then from the above results, for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$ Event $\mathcal{A}$ occurs with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ and we have
@@
To prove the consistency of our estimator $\hat{\Delta}$ we use the framework of Negahban et al. (2012) which requires showing that our penalty, $\|\ \cdot\|_{1}$, is decomposable, that our D-trace loss function satisfies the restricted strong convexity (RSC) property, and that we suitably choose a penalty $\tau_{n_{1},n_{2}}$. Specifically we will use Theorem 1 from Negahban et al. (2012).
Decomposability of $\|\cdot\|_{1}$.
The decomposability of the penalty term is defined in Definition 1 of Negahban et al. (2012). For our loss, we let $S_{\Delta}$ be the support of $\Delta$, the true difference. Let ¯=$\overline{\mathcal{M}}=\left\{\theta:\mathbb{R}^{2p\times 2p}\big{|}\theta_{ij}\neq 0\text{ for all }(i,j)\in S_{\Delta}\right\}$ and ¯=$\overline{\mathcal{M}}^{\perp}=\left\{\gamma:\mathbb{R}^{2p\times 2p}\big{|}\gamma_{ij}=0\text{ for all }(i,j)\in S_{\Delta}\right\}$. Then $\|\cdot\|_{1}$ is decomposable since $\|\theta+\gamma\|_{1}=\|\theta\|_{1}+\|\gamma\|_{1}$ for all $\theta\in\overline{\mathcal{M}}$ and $\gamma\in\overline{\mathcal{M}}^{\perp}$.
RSC of D-trace Loss.
Next we must show that the D-trace loss satisfies the RSC condition from Definition 2 of Negahban et al. (2012). We must show that $\mathbf{m}^{T}\left(\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\right)\mathbf{m}>\kappa_{L}\|\mathbf{m}\|_{2}^{2}$ for all $\mathbf{m}\in\mathcal{C}(S_{\Delta})$ where $\mathcal{C}(S_{\Delta})=\left\{\mathbf{m}\in\mathbb{R}^{4p^{2}}\big{|}\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}\leq 3\|\mathbf{m}_{S_{\Delta}}\|_{1}\right\}$ for some $\kappa_{L}$. The notation $\mathbf{m}_{S_{\Delta}}$ selects the entries of $\mathbf{m}$ that correspond to the non-zero indices of $\Delta$.
First note that for $\mathbf{m}\in\mathcal{C}(S_{\Delta})$, $\|\mathbf{m}\|_{1}\leq 4\|\mathbf{m}_{S_{\Delta}}\|_{1}\leq 4\sqrt{s_{\Delta}}\|\mathbf{m}_{S_{\Delta}}\|_{2}$ where $s_{\Delta}=|S_{\Delta}|$. Recall $\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})=0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma_{2}}\otimes\hat{\Sigma}_{1})$. Showing RSC proceeds the same as in Wang et al. (2021).
@@
Recall we showed that for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$, Event $\mathcal{A}$ occurs with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$. Using this and the fact that $\lambda_{\mathrm{min}}(\Sigma_{l})=\lambda_{\mathrm{min}}(f_{l})$ we have that
@@
Choice of Penalty $\boldsymbol{\tau_{n_{1},n_{2}}}$.
The last ingredient in our theoretical analysis is the choice of penalty scale $\tau_{n_{1},n_{2}}$. Theorem 1 in Negahban et al. (2012) we need $\tau_{n_{1},n_{2}}\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$. We know that $2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}=2\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}.$
Let $\Gamma=0.5\left(\Sigma_{2}\otimes\Sigma_{1}+\Sigma_{1}\otimes\Sigma_{2}\right)$ and $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}+\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}\right).$
Then we write
@@
Note that we showed in (6) that for $\min(n_{1},n_{2})\geq\frac{C_{1}^{3}}{C_{f}^{3}}8^{9/2}$ we have with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
$\|\hat{\Gamma}-\Gamma\|_{\infty}\leq\frac{C_{1}C_{f}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\,.$
Thus we get
@@
$\leq\frac{\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}\,.$
@@
$\tau_{n_{1},n_{2}}\geq\frac{2\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}$
@@
Convergence Rate for Difference in Networks.
Now we are ready to state the result which uses Corollary 1 of Negahban et al. (2012).
Under Assumption 2 for $\tau_{n_{1},n_{2}}\geq\frac{2}{\min(n_{1},n_{2})^{1/3}}\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)$,
$\min(n_{1},n_{2})\geq\max\left(8^{9/2}C_{1}^{3}/C_{f}^{3},\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}\right)$, and $M_{n,l}=O(n_{l}^{2/3})$ we have with probability greater than $1-C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ for any $H>0$,
$\|\hat{\Delta}-\Delta\|_{F}\leq\frac{\sqrt{18s_{\Delta}}\tau_{n_{1},n_{2}}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}$
Since $\tau_{n_{1},n_{2}}$ is of order $\min(n_{1},n_{2})^{-1/3}$, convergence of our D-trace estimator is also of order $\min(n_{1},n_{2})^{-1/3}$.
A.4 Simulation Specifics and Results

### v2

@@ Appendix B FUNCTIONAL DEPENDENCE FRAMEWORK
Appendix C THEORETICAL RESULTS
To prove Theorem 1 we will use the framework of Negahban et al. (2012), specifically Corollary 1. This requires establishing that the D-trace loss satisfies the Restricted Strong Convexity (RSC) condition and that the regularizing penalty is decomposable. An essential ingredient to all these results is a concentration bound on $\hat{f}_{l}$. We begin by establishing this bound.
Lemma 1 (Concentration inequality on $\hat{f}_{l}$).
Under Assumption 1, for any $H>0$, and for an optimal bandwidth choice of $M_{n_{l}}^{*}=O(n_{l}^{2/3})$, there exists a constant $C_{1}>0$ that depends only on $\kappa$ and $C_{0}$ such that
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,,$ (3)
where $l\in\{1,2\}$ indicates condition and the constant $C_{2}>0$ is constant in $n_{l}$ but depends on $C_{1}$ and $H$.
Proof.
Under Assumption 1 we have that Theorem 3.1 from (Fiecas et al., 2019) holds so that for any $\delta>0$ and $H>0$ and for $C_{1}>0$ depending only on $\kappa$ and $C_{0}$
$\mathbb{P}\left(\sup_{\{\lambda_{j}=2\pi j/n_{l},\lfloor(n_{l}-1)/2\rfloor\leq j\leq\lfloor n_{l}/2\rfloor\}}\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}>\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right)\leq C_{2}p^{2}n_{l}^{-H}\,,$
where $l$ indexes condition 1 or 2 and $C_{2}$ is constant in $n_{l}$ but depends on $C_{1}$ and $H$. Note that this same inequality holds for both conditions $l=1,2$ because Assumption 1 holds for both conditions. We allow $n_{l}$ to differ between conditions for full generality. We can then select $M_{n_{l}}$ to minimize the deviations. That is we select $M_{n_{l}}$ to minimize $\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}$. The algebra is omitted but setting $\frac{\partial{}}{\partial{M_{n_{l}}}}\left[\frac{C_{1}M_{n_{l}}}{n_{l}}+\frac{8n_{l}^{\delta}}{M_{n_{l}}^{1/2+\delta}}\right]=0$ and solving we get $M^{*}_{n_{l}}=\left(\frac{8}{C_{1}}\left(\frac{1}{2}+\delta\right)\right)^{2/(3+2\delta)}n_{l}^{\frac{2+2\delta}{3+2\delta}}$. It can be shown that the second derivative is positive on $\delta>0$ so this is indeed a global minimum (on $\delta>0$). Plugging this back into the equation yields
@@ Appendix C THEORETICAL RESULTS
Letting $\delta\rightarrow 0$ the RHS of our concentration bound becomes $C_{1}8^{2/3}2n_{l}^{-\frac{1}{3}}$ establishing the lemma.
Remark 1.
Since $\|\hat{\Sigma}_{l}(\lambda_{j})-\Sigma_{l}(\lambda_{j})\|_{\infty}\leq\|\hat{f}_{l}(\lambda_{j})-f_{l}(\lambda_{j})\|_{\infty}$, it follows that Lemma 1 also holds for $\|\hat{\Sigma}_{l}(\lambda_{j})-\Sigma_{l}(\lambda_{j})\|_{\infty}$.
To establish RSC of the D-trace loss function we will also need a concentration inequality on the second derivative of the D-trace loss which is stated in Lemma 2.
Lemma 2 (Concentration inequality on second derivative of D-trace loss).
Under the conditions from Lemma 1 we have that for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$, with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
$\left\|0.5\left(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}\right)-0.5\left(\Sigma_{1}\otimes\Sigma_{2}+\Sigma_{2}\otimes\Sigma_{1}\right)\right\|_{\infty}\leq\frac{C_{f}C_{1}8^{5/2}}{\min(n_{1},n_{2})^{1/3}}\,,$
where $C_{f}=\max(\|f_{1}\|_{\infty},\|f_{2}\|_{\infty})$ and $C_{1}$ is the same as in Lemma 1.
Proof.
Begin by noting that both
@@
Also we have that $\|\Sigma_{l}\|_{\infty}\leq\|f_{l}\|_{\infty}$. Using Lemma 1, we get that with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
@@
This also holds for $\|\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}-\Sigma_{2}\otimes\Sigma_{1}\|_{\infty}$. Defining
@@
then from the above results, for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$, Event $\mathcal{A}$ occurs with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ and we have
@@
Using Lemma 1 and Lemma 2 we can establish that the D-trace loss satisfies the restricted strong convexity condition. This is formalized in Lemma 3
Lemma 3 (RSC of D-trace loss).
Under the conditions of Lemma 1, for $\min(n_{1},n_{2})\geq\max\left(8^{9/2}C_{1}^{3}/C_{f}^{3},\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}\right)$ with probability ¿ $1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$ RSC holds for the D-trace loss with $\kappa_{L}=\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})/2$. We use $\lambda_{\mathrm{min}}(A)$ to denote the minimum eigenvalue of a matrix $A$ and $s_{\Delta}$ to denote the number of non-zero entries of $\Delta$. The constant $C_{f}$ is defined in Lemma 2 while $C_{1},C_{2}$ are defined in Lemma 1.
Proof.
To show that the D-trace loss function satisfies the RSC condition, we must show that $\mathbf{m}^{T}\left(\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\right)\mathbf{m}>\kappa_{L}\|\mathbf{m}\|_{2}^{2}$ for all $\mathbf{m}\in\mathcal{C}(S_{\Delta})$ where $\mathcal{C}(S_{\Delta})=\left\{\mathbf{m}\in\mathbb{R}^{4p^{2}}\big{|}\|\mathbf{m}_{S_{\Delta}^{c}}\|_{1}\leq 3\|\mathbf{m}_{S_{\Delta}}\|_{1}\right\}$ for some $\kappa_{L}$. For our loss, we let $S_{\Delta}$ be the support of $\Delta$, the true difference. The notation $\mathbf{m}_{S_{\Delta}}$ selects the entries of $\mathbf{m}$ that correspond to the non-zero indices of $\Delta$ while $\mathbf{m}_{S_{\Delta}^{c}}$ selects the entries corresponding to the zero indices of $\Delta$.
First note that for $\mathbf{m}\in\mathcal{C}(S_{\Delta})$, $\|\mathbf{m}\|_{1}\leq 4\|\mathbf{m}_{S_{\Delta}}\|_{1}\leq 4\sqrt{s_{\Delta}}\|\mathbf{m}_{S_{\Delta}}\|_{2}$. Recall $\nabla^{2}L(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})=0.5(\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}+\hat{\Sigma_{2}}\otimes\hat{\Sigma}_{1})$. Showing RSC proceeds the same as in Wang et al. (2021).
@@
Using Lemma 2 and the fact that $\lambda_{\mathrm{min}}(\Sigma_{l})=\lambda_{\mathrm{min}}(f_{l})$ we have that, for $\min(n_{1},n_{2})\geq 8^{9/2}C_{1}^{3}/C_{f}^{3}$ and with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
@@
We are now ready to prove Theorem 1.
Proof of Theorem 1.
We first proceed by establishing the decomposability of our penalty term, $\left\|\Delta\right\|_{1}$. Decomposability is defined in Definition 1 of Negahban et al. (2012). Let ¯=$\overline{\mathcal{M}}=\left\{\theta:\mathbb{R}^{2p\times 2p}\big{|}\theta_{ij}\neq 0\text{ for all }(i,j)\in S_{\Delta}\right\}$ and ¯=$\overline{\mathcal{M}}^{\perp}=\left\{\gamma:\mathbb{R}^{2p\times 2p}\big{|}\gamma_{ij}=0\text{ for all }(i,j)\in S_{\Delta}\right\}$. Then $\|\cdot\|_{1}$ is decomposable since $\|\theta+\gamma\|_{1}=\|\theta\|_{1}+\|\gamma\|_{1}$ for all $\theta\in\overline{\mathcal{M}}$ and $\gamma\in\overline{\mathcal{M}}^{\perp}$.
Next we discuss a suitable choice of penalty scale $\tau_{n_{1},n_{2}}$. From Theorem 1 in Negahban et al. (2012) we require $\tau_{n_{1},n_{2}}\geq 2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}$. We know that $2\|\nabla L_{D}(\Delta,\hat{\Sigma}_{1},\hat{\Sigma}_{2})\|_{\infty}=2\|0.5\left(\hat{\Sigma}_{1}\Delta\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\Delta\hat{\Sigma}_{1}\right)-\left(\hat{\Sigma}_{1}-\hat{\Sigma}_{2}\right)\|_{\infty}.$
Let $\Gamma=0.5\left(\Sigma_{2}\otimes\Sigma_{1}+\Sigma_{1}\otimes\Sigma_{2}\right)$ and $\hat{\Gamma}=0.5\left(\hat{\Sigma}_{2}\otimes\hat{\Sigma}_{1}+\hat{\Sigma}_{1}\otimes\hat{\Sigma}_{2}\right).$ Then we write
@@
By Lemma 2 we have for $\min(n_{1},n_{2})\geq\frac{C_{1}^{3}}{C_{f}^{3}}8^{9/2}$, with probability $>1-2C_{2}p^{2}\min(n_{1},n_{2})^{-H}$,
@@
$\leq\frac{\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}\,.$ (6)
@@
$\tau_{n_{1},n_{2}}\geq\frac{2\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)}{\min(n_{1},n_{2})^{1/3}}\,,$
@@
Combining the above with Lemma 3 we get that for $\tau_{n_{1},n_{2}}\geq\frac{2}{\min(n_{1},n_{2})^{1/3}}\left(C_{1}C_{f}8^{5/2}\|\Delta\|_{1}+4C_{1}8^{3/2}\right)$, $\min(n_{1},n_{2})\geq\max\left(8^{9/2}C_{1}^{3}/C_{f}^{3},\left(\frac{32s_{\Delta}C_{f}C_{1}8^{5/2}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\right)^{3}\right)$, and $M_{n,l}=O(n_{l}^{2/3})$ we have with probability greater than $1-Cp^{2}\min(n_{1},n_{2})^{-H}$ for any $H>0$,
$\|\hat{\Delta}-\Delta\|_{F}\leq\frac{\sqrt{18s_{\Delta}}\tau_{n_{1},n_{2}}}{\lambda_{\mathrm{min}}(f_{1})\lambda_{\mathrm{min}}(f_{2})}\,,$
where we have defined $C=2C_{2}$. Since $\tau_{n_{1},n_{2}}$ is of order $\min(n_{1},n_{2})^{-1/3}$, convergence of our D-trace estimator is also of order $\min(n_{1},n_{2})^{-1/3}$.
Appendix D SIMULATION SPECIFICS AND RESULTS